# Gesture Generation (Still) Needs Improved Human Evaluation Practices: Insights from a Community-Driven State-of-the-Art Benchmark

### v1

@@ 3 Key Limitations of Current Human-Evaluation Practices
Year Venue Another model Training dataset Modelling Goal Directly compared to…
Quality Alignment Mocap Another model in the survey
2023 CVPR DiffGesture [90] TED, TED-Expressive Na, Sm Rh Another model in the survey
QPGesture [79] BEAT Hu Gen Another model in the survey
RACER [69] Trinity, own Re Gen, Sem Another model in the survey
TalkSHOW [82] SHOW Gen Another model in the survey
SG. Bodyformer [60] Trinity, TWH Hu Gen Another model in the survey
GestureDiffuCLIP [5] BEAT, ZEGGS Hu Sem, St Another model in the survey
LDA [4] Trinity, ZEGGS Pref St Another model in the survey
ICCV C-DiffGAN [2] PATS Na Sem, Rh, St Another model in the survey
LivelySpeaker [89] BEAT, TED Na, Sm Sem Another model in the survey
2024 CVPR AMUSE [20] BEAT2 Rh, Em Another model in the survey
Audio2Photoreal [58] Audio2Photoreal Re Another model in the survey
ConvoFusion [55] DnD Na Gen, Sem Another model in the survey
DiffSHEG [16] BEAT, SHOW Re Rh Another model in the survey
EMAGE [47] BEAT2 Re Another model in the survey
EmoTransition [65] BEAT-ETrans, TED-ETrans Na, Sm Another model in the survey
ProbTalk [52] SHOW Another model in the survey
SG. Sem. Gest. [87] BEAT, ZEGGS, own Hu Sem, Rh Another model in the survey
SG. Asia SIGGesture [18] BEAT Na Sem, Rh Another model in the survey
2025 CVPR HOP [17] TED, TED-Expressive Na, Sm Sem, Rh Another model in the survey
LOM [15] BEAT2 Another model in the survey
RAG-Gesture [56] BEAT2 Na Sem Another model in the survey
SG. MeCo [14] BEAT2, ZEGGS Hu Gen Another model in the survey
ICCV GestureHydra [77] SHOW, own Na Sem Another model in the survey
GestureLSM [49] BEAT2 Sm, Re Rh Another model in the survey
SemGes [48] BEAT, TED-Expressive Na Sem, Rh Another model in the survey
SemTalk [86] BEAT2, SHOW Re Sem, Rh Another model in the survey

### v2

# Towards Reliable Human Evaluations in Gesture Generation: Insights from a Community-Driven State-of-the-Art Benchmark

@@ 3 Key Limitations of Current Human-Evaluation Practices
Year Venue Model Training dataset Modelling Goal Directly compared to…
Quality Alignment Mocap A model in the survey
2023 CVPR DiffGesture [92] TED [85], TED-Expr. [92] Na, Sm Rh A model in the survey
QPGesture [81] BEAT [48] Hu Gen A model in the survey
RACER [71] Trinity [30], own Re Gen, Sem A model in the survey
TalkSHOW [84] SHOW [84] Gen A model in the survey
SG. Bodyformer [62] Trinity [30], TWH [46] Hu Gen A model in the survey
GestureDiffuCLIP [6] BEAT [48], ZEGGS [31] Hu Sem, St A model in the survey
LDA [5] Trinity [30], ZEGGS [31] Pref St A model in the survey
ICCV C-DiffGAN [3] PATS [2] Na Sem, Rh, St A model in the survey
LivelySpeaker [91] BEAT [48], TED [85] Na, Sm Sem A model in the survey
2024 CVPR AMUSE [21] BEAT2 [49] Rh, Em A model in the survey
Audio2Photoreal [60] own Re A model in the survey
ConvoFusion [57] DnD [57] Na Gen, Sem A model in the survey
DiffSHEG [17] BEAT [48], SHOW [84] Re Rh A model in the survey
EMAGE [49] BEAT2 [49] Re A model in the survey
EmoTransition [67] own Na, Sm A model in the survey
ProbTalk [54] SHOW [84] A model in the survey
SG. Sem. Gest. [89] BEAT [48], ZEGGS [31], own Hu Sem, Rh A model in the survey
SG. Asia SIGGesture [19] BEAT [48] Na Sem, Rh A model in the survey
2025 CVPR HOP [18] TED [85], TED-Expr. [92] Na, Sm Sem, Rh A model in the survey
LOM [16] BEAT2 [49] A model in the survey
RAG-Gesture [58] BEAT2 [49] Na Sem A model in the survey
SG. MeCo [15] BEAT2 [49], ZEGGS [31] Hu Gen A model in the survey
ICCV GestureHydra [79] SHOW [84], own Na Sem A model in the survey
GestureLSM [51] BEAT2 [49] Sm, Re Rh A model in the survey
SemGes [50] BEAT [48], TED-Expr. [92] Na Sem, Rh A model in the survey
SemTalk [88] BEAT2 [49], SHOW [84] Re Sem, Rh A model in the survey